OpenStudy (texaschic101):

can someone help me set this up please...the wording is kinda confusing me :)

OpenStudy (texaschic101):

Half of the largest of 3 consecutive integers is 1/4 of 24 less then thrice the second integer. What are the numbers. So I have x, x+1 , x + 2 then I have : 1/2(x + 2) = 1/4(3(x + 1) - 24.....Is this correct ?

OpenStudy (texaschic101):

@amistre64 ....please help

OpenStudy (texaschic101):

@ash2326 ...please take a look at this

OpenStudy (texaschic101):

or is it 1/4 of 24, being 24/4 = 6 3(x + 1) - 6 ???

OpenStudy (texaschic101):

@satellite73 ...help me please

OpenStudy (texaschic101):

@ivettef365 ....help

OpenStudy (amistre64):

Half of the largest of 3 consecutive integers is 1/4 of 24 less then thrice the second integer a, b, c c/2 = (3b-24)/4

OpenStudy (texaschic101):

so I would multiply the equation by 4...right

OpenStudy (amistre64):

sure ... you might want to consider x-2, x-1, x as the numbers since it jsut uses x and x-1

OpenStudy (texaschic101):

2c = 3b - 24 wait...I am confused with the c and b.....is c supposed to be x + 2 and b is x + 1 ?

OpenStudy (amistre64):

a,b,c are the integers for simplicity .... let them be: x-2, x-1, x so that you done have to do extra workings

OpenStudy (amistre64):

2x = 3(x-1)-24

OpenStudy (texaschic101):

2x = 3x + 3 - 24 2x - 3x = - 21 -x = -21 x = 21 x - 1 = 20 x - 2 = 19 ???

OpenStudy (amistre64):

watch your signs ...

OpenStudy (texaschic101):

are they supposed to be x -1 and x - 2 or x + 1 and x + 2 ?

OpenStudy (amistre64):

you decide ... they could just as well be: x-1,x,x+1 whatever is more convenient for you ....

OpenStudy (amistre64):

a,b,c are the numbers; you define them in x terms

OpenStudy (texaschic101):

so x is 21...right ?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

the numbers are: a,b,c the relationship is: 2c = 3b - 24 you define a,b,c in whatever x terms make you happiest and solve

OpenStudy (amistre64):

if b=x, c=x+1 if b=x+1, c=x+2 if b=x-1, c=x they all represent the same thing; so pick one and work it thru

OpenStudy (texaschic101):

2c = 3b - 24 2(x + 1) = 3x - 24 2x + 2 = 3x - 24 2x - 3x = -24 - 2 -x = -26 x = 26 x + 1 = 27 is the third number x - 1 ?

OpenStudy (amistre64):

a,26,27 ... a must be 25

OpenStudy (amistre64):

so yes, in that arrangement; a = x-1

OpenStudy (texaschic101):

thank you soooo much...your the best

OpenStudy (amistre64):

lol, your welcome